Angle C and angle D are complementary angles, the measure of angle C happens to be 8 times as large as the measure of angle D. What are the measures of the angles?

Answer by ikleyn(52864)

.

    D + 8D = 90

    9D = 90

     D = 90/9 = 10.


ANSWER.  D = 10 degrees;   C = 8*10 = 80 degrees.